# Interpretable Responsibility Sharing as a Heuristic for Task and Motion Planning

### v1

@@ 5.3.1 Performance Comparison
ORS (ours) 0.963 $\pm$ 0.004 $0.954\pm 0.004$ 0.980 $\pm$ 0.002 0.980$\pm$0.003 11.05 $\pm$ 4.79
Table 1: Quantitative results on various household tasks. ’-’ indicates metrics do not apply to those specific baselines.
@@
We compare the performance of IRS as a heuristic for TAMP with a well-established formulation, LGP. To validate that Responsibility Sharing adapts to the environment and changes depending on the initial and goal states, we also implemented a control baseline that always utilizes auxiliary objects available in the environment. In the five-fold cross-validation test environments, IRS achieved a minimum effort of $11.05\pm 4.79$, while LGP achieved $13.47\pm 6.21$ and the control achieved $13.37\pm 5.83$ (Table 1). LGP performs better when there are fewer target objects (1 to 2) in the environment, while the control performs better with an increased number of target objects (greater than 2), showing the exponential improvement of responsibility sharing. Effort-based evaluations validate that IRS improves the agent’s effectiveness due to its dynamic heuristic selection. Additionally, the difference between the control and LGP suggests that always using auxiliary objects can improve performance but may not be the optimal behavior for the agent.
@@
Furthermore, we conducted extensive evaluations on the decision-making capacity of our proposed method, ORS. The quantitative results are shown in Table 1. ORS outperforms all baselines across three tasks, achieving significant decision-making accuracy, which is reflected in its performance as it requires the minimum effort to complete these tasks. The ANN, RRL, and CARL models show notable performance, with ANN achieving an accuracy of $80.5\pm 6.9\%$, RRL achieving an accuracy of $78.1\pm 7.7\%$, and CARL achieving a superior accuracy of $95.0\pm 5.6\%$. These accuracies highlight the effectiveness of rule-based approaches in decision-making tasks. However, our proposed method, ORS, significantly outperforms all these models, achieving an accuracy of $96.3\pm 0.4\%$. This demonstrates the robustness and reliability of ORS in decision-making tasks, achieving the highest performance metrics compared to other baseline models. While there is a clear trade-off between decision-making accuracy and interpretability among the spectrum of baselines, ORS achieves both high accuracy and interpretability. This trade-off is evident in the physical effort required by each baseline to complete the test tasks, with ORS completing the tasks with minimum effort. These results indicate that our IRS, enhanced with ORS, is capable of effectively guiding high-level planning in complex tasks.
@@ 5.3.2 Confidence Comparison
RRL $negative$longleftarrow $obj2$ $on$ $table_{east}$ |||| $tray$ $on$ $table_{north}$ |||| $jug$ $on$ $table_{west}$ 0.4188
$positive$longleftarrow $obj0$ $on$ $table_{south}$ |||| $obj1$ $on$ $table_{south}$ |||| $\neg goal$ $at$ $table_{south}$ |||| $\neg goal$ $at$ $table_{east}$ |||| $\neg action_{fill}$ 0.5938
CARL $negative$longleftarrow $obj4$ $on$ $table_{south}$ |||| $\neg goal$ $at$ $table_{handover}$ |||| $\neg obj0$ $on$ $table_{south}$ 0.6244
$positive$longleftarrow $obj0$ $on$ $table_{east}$\&& $\neg obj0$ $on$ $table_{south}$ 0.9408
ORS $negative$longleftarrow $obj0$ $on$ $table_{south}$\&& $obj1$ $on$ $table_{south}$\&& $obj2$ $on$ $table_{south}$\&& $obj3$ $on$ $table_{south}$\&& $\neg goal$ $at$ $table_{south}$ 0.7705
$positive$longleftarrow $obj0$ $on$ $table_{south}$\&& $is$ $helper$ $exist$\&& $\neg goal$ $at$ $table_{south}$\&& $\neg goal$ $at$ $table_{east}$\&& $\neg action_{fill}$\&& $\neg action_{pour}$ 0.9801
Table 2: Comparison of rule-based confidence scores across RRL, CARL, and ORS models, highlighting the varied decision-making criteria in task and motion planning. Positive rules indicate appropriate conditions for usage, while negative rules indicate otherwise.
Table 2 provides a comparative analysis of rule-based confidence scores across three models: RRL, CARL, and ORS. Each model employs distinct rules for decision-making in task and motion planning. The RRL model categorizes its rules into negative and positive based on object placements and interactions, while the CARL model uses conditions involving object interactions and goal locations, resulting in varying confidence levels for negative and positive rules. The ORS model, aimed at optimized rule synthesis, augment conditions by combining object locations and actions, yielding high confidence scores, especially for positive scenarios. This table highlights the complex interplay between spatial configurations and model predictions, showcasing the models’ capabilities to synthesize and apply rules in intricate task environments. ORS was able to generate rules with the highest confidence and interpretability, involving a higher number of rule combinations. Qualitatively, it is also observed that despite having high confidence, CARL generated less meaningful rules for the negative class. This is because there is no instance of the goal position being at the handover table, which adds no additional information to the rule combinations.
@@ 5.3.3 Responsibility Sharing in Humans
Table 3: Comparative analysis of ORS and human behavior in task execution.
We conducted human experiments to compare ORS with human subjects regarding their choices in utilizing auxiliary objects and sharing task responsibilities. These experiments focused on the serving task under constrained scenarios, where a varying number of objects and a tray were initially located on a table representing the counter. The goal positions varied among three options to create differences in both the distance (Close, Medium, Far) and the number of objects to be carried. Table 3 compares human behavior with ORS’s object utilization, providing insights into Responsibility Sharing (RS) in humans. The percentages indicate the proportion of participants whose tray usage aligned with ORS’s. The experimental setup in the simulation closely matched real-world scenarios in terms of both position and distance, enabling us to observe implicit human behavior regarding the use of auxiliary objects. Reflecting on the findings in Table 3, we emphasize that in 13 out of 15 experimental settings, ORS’s behavior matched human actions in using the tray. This highlights two points: the presence of the responsibility-sharing concept in humans and the high matching ratio indicating that ORS’s approach was human-like, even though this was not the initial intent.
@@ 5.4 Ablation Studies on Interpretability and Accuracy Trade-Off
Figure 6: Interpretability and Accuracy Trade-Off in ORS
@@ 6 Discussions
6.1 Human Bias and Responsibility Sharing
IRS is a promising framework that leverages human bias to enhance the classical TAMP formulation. As shown in Table 1, its application to various household tasks has been notably successful. IRS outperformed classical TAMP approaches, reinforcing our initial hypothesis that humans are a common factor in both environment construction and task design; therefore, an agent should utilize the auxiliary objects available in the environment. Furthermore, we investigated the practicality and effectiveness of always using an object with our control agent. The results show that solutions to sequential decision-making and manipulation problems are highly dependent on the initial configuration. This underscores the need for an adaptive use of the responsibility-sharing heuristic, which we address with ORS. Finally, our human experiments demonstrated strong parallels with the concept of responsibility sharing and human decision-making, as the embodied agent’s behavior closely matched human behavior in terms of the usage of auxiliary objects. This suggests that humans also leverage their own bias in these tasks through the use of auxiliary objects.
@@ 7 Limitations
IRS proposes a novel approach to improve the effectiveness and interpretability of existing TAMP formulations through the concept of responsibility sharing, centered around its novel decision-making mechanism, ORS. This approach brings its own challenges. ORS operates under a discrete state representation, limiting the applicability of continuous features common in robotic settings. Ideally, each auxiliary object should be represented with continuous affordance properties, such as the capacity of a tray; however, our model currently assumes that auxiliary objects can interact with all remaining objects without capacity concerns. For example, we assume a tray can carry all objects or a pitcher can fill every glass without needing to be refilled. Additionally, as mentioned in Section 4.1, our model operates in fully observable scenarios due to the absence of scene graph generation, which is beyond the scope of this study. If the initial state and goal can be observed from a scene and transformed into a first-order logic representation, then the underlying IRS mechanism is capable enough to operate under partially observable domains, assuming the TAMP formulation is supported with enough information to calculate feasible motions. In future work, we aim to incorporate continuous features into our framework to handle complex object properties and environmental settings and use scene graph generation to extend the application of IRS to partially observable settings.
@@ 8 Conclusion
We present IRS, a novel heuristic for task and motion planning that leverages human bias by incorporating responsibility sharing with auxiliary objects through its interpretable decision-making mechanism, ORS. ORS is trained with a dataset created by CPG, containing samples that indicate whether using auxiliary objects will improve the agent’s performance. The optimal integration of rules by ORS ensures transparency and applicability, making it an effective decision-making mechanism for IRS heuristic adaptation. Experiments in household tasks demonstrate that IRS significantly improves agents’ effectiveness compared to previous TAMP formulations, and ORS outperforms existing baselines, achieving the highest accuracy and interpretability.

### v2

@@ 5.3.1 Performance Comparison
[…]
Table 1: Quantitative Performance Comparison. Evaluation of IRS (using ORS) against baselines across Serving, Pouring, and Handover tasks. Metrics include decision-making performance (Accuracy, Precision, Recall, F1) and physical task efficiency (Effort, measured in joint displacement). Values represent mean $\pm$ standard deviation across 5-fold cross-validation. Bold indicates the best performance. Note that LGP and Control are fixed strategies (standard search vs. forced auxiliary usage) and do not have classification metrics (’-’).
@@
We compare the performance of IRS as a heuristic for TAMP with a well-established formulation, LGP. To validate that Responsibility Sharing adapts to the environment and changes depending on the initial and goal states, we also implemented a control baseline that always utilizes auxiliary objects available in the environment.
In the five-fold cross-validation test environments, IRS achieved a minimum effort of $11.05\pm 4.79$, while LGP achieved $13.47\pm 6.21$ and the control achieved $13.37\pm 5.83$ (Table 1). LGP performs better when there are fewer target objects (1 to 2) in the environment, while the control performs better with an increased number of target objects (greater than 2), showing the exponential improvement of responsibility sharing. Effort-based evaluations validate that IRS improves the agent’s effectiveness due to its dynamic heuristic selection. Additionally, the difference between the control and LGP suggests that always using auxiliary objects can improve performance but may not be the optimal behavior for the agent.
It is worth noting that the standard deviation for the Effort metric is relatively high across all methods. This is primarily attributed to the diversity of the test scenarios across different experimental settings, which feature a varying number of target objects (ranging from 1 to 5) and differing distances between initial and goal configurations. Consequently, the baseline physical effort required varies significantly between task instances. However, IRS achieves the lowest standard deviation ($4.79$) compared to LGP ($6.21$) and the Control ($5.83$), indicating that our method not only reduces the mean effort but also provides more consistent performance across diverse task complexities.
@@
Furthermore, we conducted extensive evaluations on the decision-making capacity of our proposed method, ORS. The quantitative results are shown in Table 1. ORS outperforms all baselines across three tasks, achieving significant decision-making accuracy, which is reflected in its performance as it requires the minimum effort to complete these tasks. The quantitative results reveal a clear performance hierarchy. The ANN, RRL, and CARL models show notable performance, with ANN achieving an accuracy of $80.5\pm 6.9\%$ and RRL achieving $78.1\pm 7.7\%$. Among the baselines, CARL demonstrates superior predictive power with an accuracy of $95.0\pm 5.6\%$, indicating that its relational rules are the primary drivers of performance.
However, our proposed method, ORS, significantly outperforms all baselines, achieving the highest accuracy of $96.3\pm 0.4\%$. This enhancement is due to the synergy between the learners: ORS synthesizes the high-performing relational rules from the joint rule pool with the structural, interpretable logic. By employing a bounded greedy search to optimize a Balance Score, ORS filters out unstable rules, reducing the standard deviation by an order of magnitude (from $\pm 5.6\%$ in CARL to $\pm 0.4\%$ in ORS). Thus, ORS leverages CARL for raw predictive power while relying on RRL’s structural constraints to ensure stability and interpretability.
Error Analysis and Geometric Variability: While ORS achieves high accuracy ($\approx 96.3\%$), the majority of the remaining misclassifications highlight the distinction between the continuous geometric costs used for ground-truth labeling and the discrete logical predicates used for inference. The CPG module generates labels based on the precise cost of executed motion plans; consequently, it effectively identifies scenarios where auxiliary objects are present but geometrically suboptimal (e.g., located at the far edge of the workspace) as negative. In contrast, the ORS model operates on high-level logical features (e.g., tray_on_tabletray\_on\_table) that abstract away specific metric coordinates. As a result, the model learns generalized semantic rules, such as utilizing a tray for multiple objects, which may occasionally contradict the specific geometric ground truth in boundary cases. These discrepancies represent a trade-off inherent to the framework, prioritizing the learning of robust, interpretable strategies over the overfitting of specific geometric initializations.
@@ 5.3.2 Confidence Comparison
RRL negativenegative $\longleftarrow$ $obj2$ onon $table_{east}$ $|$ traytray onon $table_{north}$ $|$ jugjug onon $table_{west}$ 0.4188
positivepositive $\longleftarrow$ $obj0$ onon $table_{south}$ $|$ $obj1$ onon $table_{south}$ $|$ $\neg goal$ atat $table_{south}$ $|$ $\neg goal$ atat $table_{east}$ $|$ $\neg action_{fill}$ 0.5938
CARL negativenegative $\longleftarrow$ $obj4$ onon $table_{south}$ $|$ $\neg goal$ atat $table_{handover}$ $|$ $\neg obj0$ onon $table_{south}$ 0.6244
positivepositive $\longleftarrow$ $obj0$ onon $table_{east}$ &\& $\neg obj0$ onon $table_{south}$ 0.9408
ORS negativenegative $\longleftarrow$ $obj0$ onon $table_{south}$ &\& $obj1$ onon $table_{south}$ &\& $obj2$ onon $table_{south}$ &\& $obj3$ onon $table_{south}$ &\& $\neg goal$ atat $table_{south}$ 0.7705
positivepositive $\longleftarrow$ $obj0$ onon $table_{south}$ &\& isis helperhelper existexist &\& $\neg goal$ atat $table_{south}$ &\& $\neg goal$ atat $table_{east}$ &\& $\neg action_{fill}$ &\& $\neg action_{pour}$ 0.9801
Table 2: Sample Rules and Confidence Scores. A comparison of extracted rules from RRL, CARL, and ORS. The Sample Rules column shows the logical preconditions that trigger a prediction. Positive rules indicate that utilizing an auxiliary object reduces physical effort (Responsibility Sharing), while Negative rules advise against it. ORS synthesizes rules with higher specificity (combining multiple spatial and action predicates via conjunctions &\&), yielding significantly higher confidence scores compared to the baselines. This illustrates ORS’s ability to capture complex, context-dependent dependencies.
Table 2 provides a comparative analysis of rule-based confidence scores across three models: RRL, CARL, and ORS. Each model employs distinct rules for decision-making in task and motion planning. The RRL model categorizes its rules into negative and positive based on object placements and interactions, while the CARL model uses conditions involving object interactions and goal locations, resulting in varying confidence levels for negative and positive rules.
The ORS model, aimed at optimized rule synthesis, augment conditions by combining object locations and actions, yielding high confidence scores, especially for positive scenarios. This table highlights the complex interplay between spatial configurations and model predictions, showcasing the models’ capabilities to synthesize and apply rules in intricate task environments. ORS was able to generate rules with the highest confidence and interpretability, involving a higher number of rule combinations. Qualitatively, it is also observed that despite having high confidence, CARL generated less meaningful rules for the negative class. This is because there is no instance of the goal position being at the handover table, which adds no additional information to the rule combinations.
@@ 5.3.3 Responsibility Sharing in Humans
Table 3: Alignment between Human and ORS Decision-Making. The table presents the percentage of human participants ($N=6$) whose decision to utilize the tray matched the decision predicted by the ORS model. The evaluation covers scenarios with varying object counts (1–5 mugs) and transport distances (Close, Medium, Far) between the counter and the target table. A value of 100% indicates perfect consensus between human intuition and the learned logical rule. The divergence in the 2-mug/Far case highlights a decision boundary where the utility of the tray becomes ambiguous for humans, whereas ORS maintains a consistent strategy.
We conducted human experiments to compare ORS with human subjects regarding their choices in utilizing auxiliary objects and sharing task responsibilities. These experiments focused on the serving task under constrained scenarios, where a varying number of objects and a tray were initially located on a table representing the counter. The goal positions varied among three options to create differences in both the distance (Close, Medium, Far) and the number of objects to be carried. Table 3 compares human behavior with ORS’s object utilization, providing insights into Responsibility Sharing (RS) in humans. The percentages indicate the proportion of participants whose tray usage aligned with ORS’s. The experimental setup in the simulation closely matched real-world scenarios in terms of both position and distance, enabling us to observe implicit human behavior regarding the use of auxiliary objects. Reflecting on the findings in Table 3, we emphasize that in 13 out of 15 experimental settings, ORS’s behavior matched human actions in using the tray.
This highlights two points: the presence of Human-Centric Environmental Bias in human decision-making, and the high matching ratio indicating that ORS effectively captures this intrinsic environmental structure, even though imitating human behavior was not the primary objective.
@@ 5.4 Ablation Studies on Interpretability and Accuracy Trade-Off
Figure 6: Impact of Rule Complexity on Model Performance. The graph illustrates the trade-off between predictive accuracy and interpretability within the ORS framework. The x-axis represents the number of rule components selected. The left y-axis tracks the Average Accuracy for both positive and negative rules, while the right y-axis measures Interpretability (defined as the Average Rule Length). As the number of components increases, the rules become more descriptive (higher interpretability) but less generalizable, leading to a drop in accuracy. ORS optimizes this trade-off by selecting a rule set that balances high accuracy with sufficient causal descriptiveness.
@@ 6 Discussions
6.1 Human-Centric Environmental Bias and Responsibility Sharing
IRS is a promising framework that leverages human-centric environmental bias to enhance the classical TAMP formulation. As shown in Table 1, its application to various household tasks has been notably successful. IRS outperformed classical TAMP approaches, reinforcing our initial hypothesis that humans are a common factor in both environment construction and task design; therefore, an agent should utilize the auxiliary objects available in the environment. Furthermore, we investigated the practicality and effectiveness of always using an object with our control agent. The results show that solutions to sequential decision-making and manipulation problems are highly dependent on the initial configuration. This underscores the need for an adaptive use of the responsibility-sharing heuristic, which we address with ORS. Finally, our human experiments demonstrated a consistent alignment between the concept of responsibility sharing and human decision-making. By comparing human decisions in physical trials against the ORS model’s predictions for semantically equivalent scenarios within the simulation, we verified that the choice to utilize auxiliary objects was driven by a strong, intuitive signal inherent in the environment. These results validate the existence of Human-Centric Environmental Bias, showing that humans naturally leverage these environmental structures (e.g., trays) to optimize effort. By capturing this behavior through interpretable rules rather than black-box approximations, ORS provides a transparent mechanism that aligns robotic planning with the logical heuristics used by humans.
@@ 7 Limitations
IRS proposes a novel approach to improve the effectiveness and interpretability of existing TAMP formulations through the concept of responsibility sharing. However, this approach entails specific limitations, particularly regarding physical realism.
Physical Capacity and Constraints: Currently, ORS operates under a discrete state representation that assumes auxiliary objects have sufficient capacity to handle assigned tasks (e.g., a tray can carry all target objects). We acknowledge that this is a critical assumption. However, the modular design of IRS provides a concrete path to relax this assumption in future work:
  1. 1.
    Solver-Level Integration: The underlying Mini-LGP solver (Eq. 9) can enforce capacity constraints during the sub-problem optimization phase. If a heuristic-suggested plan violates physical limits (e.g., objects do not fit on the tray), the solver will deem the sub-goal infeasible, forcing the planner to fall back to standard search or split the task into multiple trips.
  2. 2.
    Feature-Level Integration: To make the decision-making proactive rather than reactive, continuous affordance properties (e.g., $V_{\textrm{objects}}$ vs. $C_{\textrm{tray}}$, for volume and capacity, respectively) can be integrated into the ORS feature space $\Phi$. This would allow the rule synthesizer to learn capacity-aware conditions (e.g., Use Tray IF Count >> 2 AND FitsOnTray), bridging the gap between logical heuristics and physical constraints.
Partial Observability: Our model currently assumes fully observable domains. Extending IRS to partially observable settings would require integrating a scene graph generation module to dynamically populate the first-order logic state representation $S$ from raw sensory data. Provided that the initial state can be inferred, the underlying reasoning mechanism of IRS remains applicable.
Cognitive Complexity: Future studies will explore complex, cognitively demanding tasks where human intuition may diverge from optimal planning. We assume that such divergences are not failures of the model, but rather indicators of tasks that exceed human cognitive efficiency. Identifying scenarios where automated planning outperforms human baselines could reveal critical opportunities for automation transformation in domestic environments.
@@ 8 Conclusion
We present IRS, a novel heuristic for task and motion planning that leverages human-centric environmental bias by incorporating responsibility sharing with auxiliary objects through its interpretable decision-making mechanism, ORS. ORS is trained with a dataset created by CPG, containing samples that indicate whether using auxiliary objects will improve the agent’s performance. The optimal integration of rules by ORS ensures transparency and applicability, making it an effective decision-making mechanism for IRS heuristic adaptation. Experiments in household tasks demonstrate that IRS significantly improves agents’ effectiveness compared to previous TAMP formulations, and ORS outperforms existing baselines, achieving the highest accuracy and interpretability.